# Heapr: Hessian-based Efficient atomic expert Pruning in Output Space

### v1

@@ 3.3 Heapr Algorithm
1. Shared Gradient Covariance Estimation. For a given expert $E_{i}$, the gradients of the loss with respect to all its constituent atomic experts’ output are identical. Therefore, rather than performing redundant computations, we execute a single backward pass to obtain the gradient for the expert’s output, $\mathbf{g}_{E_{i}}=\partial\ell/\partial E_{i}$. This shared gradient is used to compute a gradient covariance matrix $\bar{\bm{G}}_{i}$, for all atomic experts belonging to $E_{i}$, accumulated over the subset of tokens $\mathcal{T}_{i}\subseteq\mathcal{D}$ routed to $E_{i}$:
@@
This approach relies solely on standard forward and backward computations, making it both exceptionally time- and memory-efficient. The space complexity of each gradient covariance matrix is only $\mathcal{O}(d_{\text{model}}^{2})$, significantly alleviating the storage bottleneck. After computing the importance $\bar{s}_{k}$} across all micro-experts in the model, we perform a global ranking and prune the lowest $r\%$ of experts across all MoE layers. The complete and optimized procedure is summarized in Algorithm 1.
Algorithm 1 Heapr: Hessian-based Efficient Atomic Expert Pruning
@@
6:  Compute $\bar{\bm{G}}_{i}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\mathbf{g}_{E_{i}}(\mathbf{x})\mathbf{g}_{E_{i}}(\mathbf{x})^{\top}$
@@
9:  Compute $\bar{s}_{k}=\tfrac{1}{|\mathcal{T}_{k}|}\sum_{\mathbf{x}\in\mathcal{T}_{k}}\tfrac{1}{2}\,\mathbf{e}_{k}(\mathbf{x})^{\top}\bar{\bm{G}}_{i}\mathbf{e}_{k}(\mathbf{x})$
@@
4 experiment
Table 1: Performance of Heapr with DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-chat, Qwen2-57B-A14B and Qwen3-30B-A3B on seven zero-shot tasks, reported in terms of accuracy.
Ratio Method Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
@@
0% Original 0.32 0.76 0.71 0.58 0.45 0.79 0.32 0.56
20% NAEE 0.32 0.71 0.66 0.55 0.40 0.77 0.29 0.53
MoE-I2 0.26 0.71 0.68 0.49 0.38 0.73 0.29 0.50
$D^{2}$-MoE 0.30 0.74 0.69 0.55 0.41 0.76 0.31 0.54
Heapr 0.32 0.76 0.71 0.57 0.45 0.79 0.32 0.56
40% NAEE 0.23 0.67 0.67 0.41 0.32 0.69 0.26 0.46
MoE-I2 0.23 0.64 0.66 0.41 0.31 0.68 0.26 0.45
$D^{2}$-MoE 0.26 0.69 0.65 0.45 0.36 0.72 0.28 0.49
Heapr 0.30 0.74 0.69 0.52 0.41 0.76 0.30 0.53
@@
0% Original 0.31 0.70 0.66 0.59 0.40 0.79 0.35 0.54
25% MC-SMoE 0.25 0.65 0.65 0.53 0.37 - - -
HC-SMoE 0.27 0.66 0.63 0.55 0.35 0.76 0.29 0.50
Sub-MoE 0.30 0.69 0.66 0.56 0.37 - - -
Heapr 0.32 0.69 0.67 0.56 0.38 0.76 0.35 0.53
50% MC-SMoE 0.18 0.33 0.52 0.29 0.19 - - -
HC-SMoE 0.23 0.61 0.65 0.47 0.35 0.58 0.23 0.45
Sub-MoE 0.25 0.58 0.58 0.46 0.25 - - -
Heapr 0.27 0.64 0.64 0.46 0.33 0.71 0.33 0.48
@@
0% Original 0.34 0.79 0.71 0.60 0.54 0.79 0.59 0.62
25% HC-SMoE 0.22 0.64 0.61 0.40 0.35 0.59 0.41 0.46
Sub-MoE 0.25 0.70 0.66 0.47 0.44 - - -
Heapr 0.33 0.77 0.70 0.55 0.49 0.78 0.50 0.59
50% HC-SMoE 0.13 0.44 0.50 0.29 0.23 0.44 0.32 0.34
Sub-MoE 0.23 0.68 0.63 0.41 0.40 - - -
Heapr 0.25 0.67 0.63 0.38 0.41 0.67 0.36 0.48
@@
0% Original 0.33 0.75 0.74 0.63 0.46 0.81 0.39 0.59
40% NAEE 0.31 0.73 0.73 0.55 0.46 0.76 0.36 0.55
MoE-I2 0.26 0.70 0.46 0.71 0.41 0.75 0.30 0.51
$D^{2}$-MoE 0.33 0.75 0.75 0.61 0.45 0.79 0.36 0.58
Heapr 0.33 0.75 0.74 0.64 0.46 0.81 0.39 0.59
@@ 4.1 Experimental Setup
We evaluate our approach on a broad spectrum of model architectures and scales to assess its generality and effectiveness, including DeepseekMoE-16B-Base (Dai et al., 2024), Qwen1.5-MoE-A2.7B-Chat (Team, 2024a), Qwen2-57B-A14B (Team, 2024b), and Qwen3-30B-A3B. All experiences are calibrated using Wikitext-2, with our method employing 128 sequences of 2048 tokens. Notably, our method introduces no additional tunable hyperparameters.
@@
For our comparisons, we evaluate six recently proposed high-performance compression methods, including expert dropping (NAEE (Lu et al., 2024), MoE-I2 (Yang et al., 2024)), expert merging (MC-SMoE (Li et al., 2024), HC-SMoE (Chen et al., 2025)), and expert decomposition (Sub-MoE (Li et al., 2025a), $D^{2}$-MoE (Gu et al., 2025)). For fairness and transparency, the reported results in Table 1 are primarily drawn from the respective original papers, ensuring consistency.
@@
We report results on seven zero-shot benchmarks using the LM-Evaluation-Harness (version 0.4.7) (Gao et al., 2024), including HellaSwag (Zellers et al., 2019), Mathqa (Amini et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Boratko et al., 2018). These tasks collectively enable repeated and consistent evaluation of our method across varied domains and reasoning tasks.
@@ 4.2 Main Results
As shown in Table 1, Heapr achieves exceptional performance across various MoE models and compression ratios. Notably, our method delivers near-lossless compression. At pruning ratios of $20\%\sim 25\%$, Heapr matches the performance of the original models on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat. More impressively, on Qwen2-57B-A14B, Heapr maintains performance almost identical to the original model even at a high $40\%$ compression ratio. In contrast, competing new methods such as Sub-MoE, $D^{2}$-MoE, and NAEE show significant accuracy degradation under the same conditions. Furthermore, on the latest Qwen3-30B-A3B model, Heapr incurs only a minimal performance loss at a $25\%$ pruning ratio, with the average accuracy dropping slightly from $0.62$ to just $0.59$. These results strongly highlight the unique advantage of Heapr in pruning at the atomic expert level, enabling substantial model efficiency improvements while maintaining and preserving core model performance effectively.
@@
Figure 2 reports the performance of Heapr on DeepSeekMoE-16B-Base using a random 128-sample subset of WikiText-2 with $2048$ tokens under different compression ratios, where the ratio denotes the fraction of parameters removed relative to the full model size. For compression ratios below $0.4$, the pruned models retain over $93\%$ of baseline accuracy while already reducing FLOPs by $21\%$. In this regime, the accuracy curve remains nearly flat, revealing substantial redundancy among micro-experts and confirming that Heapr can effectively identify and remove them. As compression increases further, accuracy degrades gracefully, highlighting a clear trade-off between efficiency and performance. Even at an extreme compression ratio of $0.9$, the model preserves about $55\%$ of its baseline accuracy while achieving a $38\%$ reduction in FLOPs. These results demonstrate both the robustness of Heapr under moderate pruning and its effectiveness in enabling aggressive acceleration while retaining performance.
@@ 4.3 Ablations
Table 2: Comparison of layer-wise pruning (CAMERA-P and Heapr) versus global pruning (Heapr) on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat across seven zero-shot tasks, with acc_norm reported for DeepSeekMoE-16B-Base and accuracy for others.
@@
Table 3 demonstrates that the performance of our Heapr algorithm is largely unaffected by the choice of calibration data, whether they are WikiText-2 or C4 dataset. This highlights the remarkable robustness and generalizability of our method, as it consistently performs well across different calibration corpora and domains. Furthermore, the table also explores the significant impact of calibration set size on pruning performance. As the number of calibration samples increases, the model’s performance improves consistently, indicating that larger calibration sets offer richer statistical coverage, which provides more reliable and informative signals for effective compression. These results strongly suggest that our method is not only robust to variations in calibration data but also benefits from the inclusion of additional diverse samples, further enhancing its overall effectiveness and stability.
@@
In this work, we introduce Heapr, a novel method that refines expert-level pruning in MoE models by enabling a more flexible and fine-grained pruning strategy at the atomic expert level. Inspired by the principles of the Optimal Brain Surgeon (OBS) theory, we evaluate the importance of atomic experts using second-order information. By transforming the analysis from the expert parameter space to that of atomic expert parameters, and further shifting it to the atomic expert output space, we significantly reduce the computational and storage bottlenecks associated with the second-order information matrix. Heapr requires only two forward passes and one backward pass to efficiently compute the importance of atomic experts. Extensive experiments on various modern MoE models demonstrate that Heapr outperforms state-of-the-art pruning methods, achieving near-lossless pruning with pruning rates of $20\%\sim 25\%$. More importantly, our method provides a much finer-grained perspective on MoE expert pruning, which we hope will contribute to a deeper, more comprehensive understanding of MoE models. Future work will explore large-scale experiments across a wider range of model and investigate the potential of parameter compensation methods after the pruning.

### v2

@@ 3.3 Heapr Algorithm
1. Shared Gradient Covariance Estimation. For a given expert $E_{i}$, the gradients of the loss with respect to all its constituent atomic experts’ output are identical. Therefore, rather than performing redundant computations, we execute a single backward pass to obtain the gradient for the expert’s output, $\mathbf{g}_{E_{i}}=\partial\ell/\partial E_{i}$. This shared gradient is used to compute a gradient covariance matrix $\bar{\bm{G}}_{i}$ for all atomic experts belonging to $E_{i}$, accumulated over the subset of tokens $\mathcal{T}_{i}\subseteq\mathcal{D}$ routed to $E_{i}$:
@@
This approach relies solely on standard forward and backward computations, making it both exceptionally time- and memory-efficient. The space complexity of each gradient covariance matrix is only $\mathcal{O}(d_{\text{model}}^{2})$, significantly alleviating the storage bottleneck. After computing the importance $\bar{s}_{k}$} across all atomic experts in the model, we perform a global ranking and prune the lowest $r\%$ of experts across all MoE layers. The complete and optimized procedure is summarized in Algorithm 1.
Algorithm 1 Heapr: Hessian-based Efficient Atomic Expert Pruning in Output Space
@@
6:  Compute $\bar{\bm{G}}_{i}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\mathbf{g}_{E_{i}}(\mathbf{x})\mathbf{g}_{E_{i}}(\mathbf{x})^{\top}$ $\triangleright$ Space complexity $\mathcal{O}(d^{2})$
@@
9:  Compute $\bar{s}_{k}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\tfrac{1}{2}\,\mathbf{e}_{k}(\mathbf{x})^{\top}\bar{\bm{G}}_{i}\mathbf{e}_{k}(\mathbf{x})$
@@
4 Experiments
Table 1: Performance of Heapr with DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-Chat, Qwen2-57B-A14B and Qwen3-30B-A3B on seven zero-shot tasks, reported in terms of accuracy. The results marked with * are obtained from the official implementation.
Ratio Method Wiki↓ PTB↓ Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Avg.↑
@@ 4 Experiments
0% Original 6.38 9.47 0.32 0.76 0.71 0.58 0.45 0.79 0.32 0.56
20% NAEE 9.44 15.02 0.32 0.71 0.66 0.55 0.40 0.77 0.29 0.53
MoE-I2 7.69 11.59 0.26 0.71 0.68 0.49 0.38 0.73 0.29 0.50
MoE-SVD 6.92 10.48 0.31 0.75 0.70 0.53 0.42 0.76 0.31 0.54
$D^{2}$-MoE 6.84 11.10 0.30 0.74 0.69 0.55 0.41 0.76 0.31 0.54
Heapr 6.54 9.88 0.32 0.76 0.71 0.57 0.45 0.79 0.32 0.56
40% NAEE 8.55 14.47 0.23 0.67 0.67 0.41 0.32 0.69 0.26 0.46
MoE-I2 9.73 15.75 0.23 0.64 0.66 0.41 0.31 0.68 0.26 0.45
$D^{2}$-MoE 7.93 14.07 0.26 0.69 0.65 0.45 0.36 0.72 0.28 0.49
Heapr 6.80 10.86 0.30 0.74 0.69 0.52 0.41 0.76 0.30 0.53
@@
0% Original 8.12 12.97 0.31 0.70 0.66 0.59 0.40 0.79 0.35 0.54
25% MC-SMoE 12.76 17.45 0.25 0.65 0.65 0.53 0.37 - - -
HC-SMoE 11.62 16.39 0.27 0.66 0.63 0.55 0.35 0.76 0.29 0.50
Sub-MoE 9.48 14.84 0.30 0.69 0.66 0.56 0.37 - - -
Heapr 8.31 14.12 0.32 0.69 0.67 0.56 0.38 0.76 0.35 0.53
50% MC-SMoE 5e2 1e3 0.18 0.33 0.52 0.29 0.19 - - -
HC-SMoE 25.50 38.18 0.23 0.61 0.65 0.47 0.35 0.58 0.23 0.45
Sub-MoE 17.51 29.00 0.25 0.58 0.58 0.46 0.25 - - -
Heapr 9.24 17.58 0.27 0.64 0.64 0.46 0.33 0.71 0.33 0.48
@@
0% Original 8.64 15.40 0.34 0.79 0.71 0.60 0.54 0.79 0.59 0.62
25% HC-SMoE 18.86 31.11 0.22 0.64 0.61 0.40 0.35 0.59 0.41 0.46
Sub-MoE 13.59 23.48 0.25 0.70 0.66 0.47 0.44 - - -
Heapr 9.10 16.80 0.33 0.76 0.70 0.55 0.49 0.78 0.50 0.59
50% HC-SMoE 72.33 162.99 0.13 0.44 0.50 0.29 0.23 0.44 0.32 0.34
Sub-MoE 21.05 43.19 0.23 0.68 0.63 0.41 0.40 - - -
Heapr 11.22 26.29 0.25 0.67 0.63 0.38 0.41 0.67 0.36 0.48
@@
0% Original 5.12 9.18 0.33 0.75 0.74 0.63 0.46 0.81 0.39 0.59
40% NAEE 6.81 11.34 0.31 0.73 0.73 0.55 0.46 0.76 0.36 0.55
MoE-I2 24.90 77.05 0.26 0.70 0.46 0.71 0.41 0.75 0.30 0.51
$D^{2}$-MoE 8.19 11.23 0.33 0.75 0.75 0.61 0.45 0.79 0.36 0.58
Heapr 5.92 9.34 0.33 0.75 0.74 0.63 0.46 0.81 0.39 0.59
@@ 4.1 Experimental Setup
We evaluate our approach on a broad spectrum of model architectures and scales to assess its generality and effectiveness, including DeepSeekMoE-16B-Base (Dai et al., 2024), Qwen1.5-MoE-A2.7B-Chat (Team, 2024a), Qwen2-57B-A14B (Team, 2024b), and Qwen3-30B-A3B. All experiments are calibrated on WikiText-2 using 128 sequences of 2048 tokens (see Appendix B for details). Notably, our method introduces no additional tunable hyperparameters.
@@
For our comparisons, we evaluate seven recently proposed high-performance compression methods, including expert dropping (NAEE (Lu et al., 2024), MoE-I2 (Yang et al., 2024)), expert merging (MC-SMoE (Li et al., 2024), HC-SMoE (Chen et al., 2025)), and expert decomposition (Sub-MoE (Li et al., 2025a), $D^{2}$-MoE (Gu et al., 2025), MoE-SVD (Li et al., 2025b)). Baseline data were collected from prior publications, prioritizing original sources, and details are provided in Appendix B. Missing data for open-source implementations were obtained from official code.
@@
We report results on seven zero-shot benchmarks using the LM-Evaluation-Harness (version 0.4.7) (Gao et al., 2024), including HellaSwag (Zellers et al., 2019), MathQA (Amini et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Boratko et al., 2018). These tasks collectively enable repeated and consistent evaluation of our method across varied domains.
@@ 4.2 Main Results
As shown in Table 1, Heapr achieves exceptional performance across various MoE models and pruning ratios. Notably, our method delivers near-lossless compression. At pruning ratios of $20\%\sim 25\%$, Heapr matches the performance of the original models on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat. More impressively, on Qwen2-57B-A14B, Heapr maintains performance almost identical to the original model even at a high $40\%$ compression ratio. In contrast, our method outperforms recent approaches such as Sub-MoE, $D^{2}$-MoE, and NAEE under the same compression ratio. Furthermore, on the latest Qwen3-30B-A3B model, Heapr incurs only a minimal performance loss at a $25\%$ pruning ratio, with the average accuracy dropping slightly from $0.62$ to just $0.59$. These results strongly highlight the unique advantage of Heapr in pruning at the atomic expert level, enabling substantial model efficiency improvements while maintaining and preserving core model performance effectively.
@@
Figure 2 reports the performance of Heapr on DeepSeekMoE-16B-Base using a random 128-sample subset of WikiText-2 with $2048$ tokens under different compression ratios, where the ratio denotes the fraction of parameters removed relative to the full model size. For compression ratios below 0.4, the pruned models retain 93% of the baseline accuracy while achieving a 20% reduction in FLOPs. In this regime, the accuracy curve is nearly flat, indicating substantial redundancy among atomic experts and suggesting that Heapr can effectively identify and remove them. As compression increases further, accuracy degrades gracefully, highlighting a clear trade-off between efficiency and performance. Even at an extreme compression ratio of 0.9, the model retains about 38% of the baseline accuracy while achieving a 55% reduction in FLOPs. These results demonstrate the robustness of Heapr under moderate pruning and its effectiveness in enabling aggressive acceleration while retaining reasonable performance.
@@ 4.3 Ablations
Table 2: Comparison of layer-wise pruning (CAMERA-P and Heapr-L) versus global pruning (Heapr-G) on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat across seven zero-shot tasks, with acc_norm reported for DeepSeekMoE-16B-Base and accuracy for others.
@@
Impact of Pruning Granularity.
To better demonstrate the importance of atomic expert decomposition, we conduct an ablation study comparing pruning at the atomic expert level and expert level. Based on equation 8, the importance score for an expert computed via equation 13 can be expressed as the sum of the importance scores of its constituent atomic experts. As reported in Table 3, expert-level pruning behaves similarly to Expert Dropping (Lu et al., 2024): the activated experts unchanged after pruning does not lead to noticeable computational speedup. In contrast, pruning at the atomic expert level reduces the dimensionality within each expert, thereby enabling real acceleration. Empirically, atomic-level pruning consistently outperforms expert-level pruning across multiple benchmarks, highlighting its effectiveness and necessity.
Empirical Correlation of Loss and Atomic Expert Importance $s_{k}$.
Figure 3: Consistency between atomic expert normalized importance score $s_{k}$ and the change in loss. The figure plots the actual loss increase $\Delta\ell$ observed upon pruning atomic experts within 10% quantile bins (ordered by original expert index) against the cumulative importance score $s_{k}$.
In Section 2, following the principles of OBS theory, we define the atomic expert importance score $s_{k}$ for $\mathbf{e}_{\mathcal{P}}$ based on the expected change in model loss. The goal of this metric is to identify atomic experts whose removal induces the smallest increase in the overall loss. However, because both the OBS formulation and the output-space approximation neglect higher-order terms, an exact numerical match between $s_{k}$ and the empirical loss change $\Delta\ell$ is not expected. Importantly, pruning ultimately requires a reliable ranking of atomic expert importance rather than an accurate prediction of $\Delta\ell$. To evaluate the ranking quality of $s_{k}$, we infer the atomic experts on the calibration set and then group them into $10\%$ bins according to their original indices. As shown in Figure 3, the observed loss increase $\Delta\ell$ for each bin closely follows the cumulative trend of the corresponding normalized importance scores $s_{k}$. This result indicates that, despite the approximations involved, the proposed $s_{k}$ metric provides a globally consistent and reliable ranking of atomic experts. It effectively identifies atomic experts whose removal causes minimal performance degradation, thereby offering a solid basis for the Heapr algorithm and supporting the accuracy of our pruning decisions.
Table 3: Comparison of pruning granularities at the expert level and the atomic expert level, where expert importance is computed by summing the importances of its atomic experts, evaluated across seven zero-shot tasks. FLOPs $r$. denotes the FLOPs reduction ratio. Results are reported on DeepSeekMoE-16B-Base.
Ratio Level FLOPs rr.↑ Wiki↓ Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA
20% Expert 0% 6.90 31.40 75.76 71.35 57.99 44.31 78.40 30.75
Atomic Expert 8% 6.54 31.54 75.88 71.43 57.39 44.62 79.05 31.52
40% Expert 0% 8.00 30.60 73.19 63.93 51.15 42.49 77.09 28.24
Atomic Expert 21% 6.80 30.00 73.78 69.06 52.29 40.61 76.50 30.12
@@
Figure 4 shows the average accuracy with error bars over random subsets of the calibration data, indicating that the performance of our Heapr algorithm is largely unaffected by the choice of calibration data, whether they are WikiText-2 or C4 dataset. This highlights the remarkable robustness and generalizability of our method, as it consistently performs well across different calibration corpora and domains. Furthermore, the table also explores the significant impact of calibration set size on pruning performance. As the number of calibration samples increases, the model’s performance improves consistently, indicating that larger calibration sets offer richer statistical coverage, which provides more reliable and informative signals for effective compression. These results strongly suggest that our method is not only robust to variations in calibration data but also benefits from the inclusion of additional diverse samples.
@@ 5 Conclusion
In this work, we introduce Heapr, a novel method that refines expert-level pruning in MoE models by enabling a more flexible and fine-grained pruning strategy at the atomic expert level. Inspired by the principles of the Optimal Brain Surgeon theory, we evaluate the importance of atomic experts using second-order information. By transforming the analysis from the expert parameter space to that of atomic expert parameters, and further shifting it to the atomic expert output space, we significantly reduce the computational and storage bottlenecks associated with the second-order information matrix. Heapr requires only two forward passes and one backward pass to efficiently compute the importance of atomic experts. Extensive experiments on various modern MoE models demonstrate that Heapr outperforms state-of-the-art pruning methods, achieving near-lossless pruning with pruning rates of $20\%\sim 25\%$. More importantly, our method provides a much finer-grained perspective on MoE expert pruning, which we hope will contribute to a deeper, more comprehensive understanding of MoE models. Future work will explore large-scale experiments across a wider range of models and investigate the potential of parameter compensation methods after the pruning.
Acknowledgments
This work was supported in part by The National Nature Science Foundation of China (Grant NO.: 62303406), in part by Ningbo Key R&D Program (NOs.: 2025Z055, 2025Z035), in part by Yongjiang Talent Introduction Programme (Grant NO.: 2023A-194-G), in part by Zhiyuan Laboratory (NO.: ZYL2024022b).